# Understanding and Embracing Imperfection in Physical Learning Networks

### v1

@@ Appendix E Measuring Experimental Bias
For each experiment, the bias vector $\mathbf{B}$ is inferred from the gate voltage dynamics of individual tasks. For single-point tasks (all tasks in Fig. 2 and task $\beta$ in Fig. 3), we look at the long time dynamics, governed by the drift along the solution line as shown by the experiments. The rate of change of the gate voltages at long times then corresponds to projection of $\mathbf{B}$ along the solution line, $d\mathbf{G}/dt\approx\mathbf{N}\mathbf{B}$ from Eq. (21). Experimentally, the rate is not constant and to compare with the linear theory we must choose where to measure this rate. For each task, we choose a point along the solution line that is closest to a predetermined central point to the observed dynamics. These central points are [$G_{1}$,$G_{2}$] = [3.0, 3.5] V in Fig. 2 F, [3.1, 4.3] V in Fig. 2 G, and [2.5, 2.5] V in Fig. 3.
For tasks consisting in two data points (e.g. task $\alpha$ in Fig. 3), for which there is no direction of zero curvature, the bias can be completely determined by setting $d\mathbf{G}/dt=0$ in Eq. (18): $\mathbf{B}=\mathbf{M}(\mathbf{G}-\mathbf{G}^{*})$, where $\mathbf{G}$ is the stationary point obtained from the experiments. However, such measurement relies on the imprecise values of all the curvatures through the matrix $\mathbf{M}$. Instead, we consider only the projection along the smallest curvature direction (the center line depicted by the blue trajectory in the last panel of Fig. 3C).
Then, for each experiment, we find the unique value of $\mathbf{B}$ that satisfies the projections found for each of the two tasks. These values are $\mathbf{B}=(2.97,4.41)$ V/s in Fig. 1F, $(-0.16,1.98)$ V/s in Fig. 1G, and $(-1.79,0.68)$ V/s in Fig. 2.
[…]
@@ Appendix F Clamping Protocols
$\mathbf{V}_{C}\approx\mathbf{V}_{F}+\bm{\chi}\eta\delta+\mathcal{O}(\delta^{2}),$ (29)
where $\bm{\chi}=\frac{d\mathbf{V}_{C}}{dO}(O^{F})$. The learning rule then becomes
@@
$=t_{h}\left[\mathbf{B}-2A_{0}\mathbf{V}_{F}\bm{\chi}\eta\delta\right]+\mathcal{O}(\delta^{2}),$ (31)

### v2

@@ Appendix E Measuring Experimental Bias
For each experiment, the bias vector $\mathbf{B}$ is inferred from the gate voltage dynamics of individual tasks. For each single-point task of Fig. 2, we run independent experiments, one per task, to measure the bias. We look at the long time dynamics in non-periodic setups, governed by the drift along the solution line as shown in the main text. The rate of change of the gate voltages at long times then corresponds to projection of $\mathbf{B}$ along the solution line, $d\mathbf{G}/dt\approx\mathbf{N}\mathbf{B}$ from Eq. (22). Experimentally, the rate is not constant and to compare with the linear theory we must choose where to measure this rate. For each task, we choose a point along the solution line that is closest to a predetermined central point to the observed dynamics. These central points are $G_{1}$$G_{2}$] = [3.0, 3.5] V in Fig. 2 F, [3.1, 4.3] V in Fig. 2 G, and [2.5, 2.5] V in Fig. 3, see  [46].
For the two tasks in Fig. 3, we inferred the bias from the periodic trajectories. We used the same method mentioned before for the single-point task $\beta$. For task $\beta$, consisting in two data points, there is no direction of zero curvature and the bias can be completely determined by setting $d\mathbf{G}/dt=0$ in Eq. (18): $\mathbf{B}=\mathbf{M}(\mathbf{G}-\mathbf{G}^{*})$, where $\mathbf{G}$ is the stationary point obtained from the experiments. However, such measurement relies on the imprecise values of all the curvatures through the matrix $\mathbf{M}$. Instead, we consider only the projection along the smallest curvature direction (the center line depicted by the blue trajectory in the last panel of Fig. 3C).
Then, for each experiment, we find the unique value of $\mathbf{B}$ that satisfies the projections found for each of the two tasks. These values are $\mathbf{B}=(2.9,4.7)$ V/s in Fig. 2F, (0.3,2.0)(-0.3,2.0) V/s in Fig. 2G, and (1.8,0.6)(-1.8,0.6) V/s in Fig. 3.
We note one change between here and previous work, which is the removal of a drainage resistor (34 k ohms in [21] Fig 7), designed to funnel the capacitor charging current to ground when learning is turned off. This resistor biased learning updates, effectively shuttling charge between the capacitor and ground when learning is toggled. We reintroduce this resistor only in Fig. 2 C and G, and only for a single network edge ($G_{1}$), in order to explore the effect of varying bias.
@@ Appendix F Clamping Protocols
$\mathbf{V}_{C}\approx\mathbf{V}_{F}+\chi\eta\delta+\mathcal{O}(\delta^{2}),$ (31)
where $\chi=\frac{d\mathbf{V}_{C}}{dO}(O^{F})$. The learning rule then becomes
@@
$=t_{h}\left[\mathbf{B}-2A_{0}\mathbf{V}_{F}\chi\eta\delta\right]+\mathcal{O}(\delta^{2}),$ (33)